# Vejde: A Framework for Inductive Deep Reinforcement Learning Based on Factor Graph Color Refinement

### v1

@@ 3.5 Value Estimate
where
$\widetilde{Q}(u,a)=h_{u}\cdot W_{a};\ W\in\mathbb{R}^{D\times|A|},\ h_{u}\in H.$
. This formulation can be seen as dividing a single step in the MDP into two sub-steps, where the first sub-step solely consists of the agent choosing a predicate symbol. If we were to extend the number of object arguments for actions, the number of sub-steps would increase to account for the additional choices.
3.6 Implementation Details
This subsection contains details which are less important in theory, but are relevant to the practical aspects of the implementation.
3.6.1 Batching
Since each sample can vary in size, unless padding is used batching for gradient descent can not be done in the common way of stacking vectors. We opt for the batching method used by graph learning code libraries such as PyTorch Geometric (Fey & Lenssen, 2019), where the adjacency matrix of the graph is stored in a COO matrix format. Thus, the node and adjacency vectors for each graph are concatenated in order to produce a single, albeit disconnected, graph.
3.6.2 Action Masking
3.6.3 Code Library Design
Vejde is designed to be domain-agnostic, and is built upon the Gymnasium interface (Towers et al., 2024). The heart of the implementation is a relational data model class, where predicate symbols, action symbols and object types of the particular problem domain are specified. This class is then used to shape the GNN and construct the graph representation of the state. We use a RDDL-specific instance of this class for our evaluation, which pulls the required information from a RDDL domain specification. In order to apply the library to a new problem domain, a relational data model for that domain has to be defined.
@@ 4 Evaluation
We evaluate Vejde using a set of problems defined with the Relational Dynamic Influence Diagram Language (RDDL) and that are executed in the Python library PyRDDLGym [^8] https://github.com/pyrddlgym-project/pyRDDLGym/commit/f7dd1dd (Taitler et al., 2024). RDDL is a description language that can be used for specifying Markov decision problems as dynamic Bayesian networks (Sanner, 2010). We test the ability of Vejde to model inductive decision policies in two ways: supervised learning from examples and reinforcement learning.
@@
We compare Vejde policies with policies parametrized using MLPs, and a planning algorithm named Prost (Keller & Helmert, 2013), the winner of the probabilistic track of the 2014 International Probabilistic Planning Competition (IPPC(Vallati et al., 2015). Prost uses a tree-search algorithm framework to find an optimal action for each state based on a series of simulations initiated from the state. As such, for each state Prost encounters it needs to be able to run thousands of trials to construct a search-tree. We think it should be noted that this is a fundamentally different problem-solving approach than the reinforcement learning method used to train the Vejde and MLP policy functions. A policy function takes a state, which can be represented purely as data, and maps it to actions. A simulator is only used during training to find the parameters of the policy, whereas Prost always needs access to a simulator of the problem in order to perform searches. We use an existing interface for Prost to PyRDDLGym for all experiments. [^9] https://github.com/pyrddlgym-project/pyRDDLGym-prost/commit/248d5d2 For the MLP policies, we used the implementation included in the pyRDDLGym set of libraries, which in turn is based on Stable Baselines 3.101010https://github.com/pyrddlgym-project/pyRDDLGym-rl/commit/9714392 In the implementation, the state is represented as a vector over all possible grounded atoms, the length of which is dependent on the language and number of objects in the problem instance. This means that the total parameter count of the MLPs input and output layers varies between instances and domains, but all use two hidden layers with 64 latent parameters each. The MLP policies are transductive, and can not be used for multiple problem instances due to the varying input and output shapes. We thus train one MLP agent for each domain and instance combination. Including another inductive agent for comparison would have been useful, but we were unable to find a single method that cover the set of problem domains as we are evaluating on. We should note, however, that there is overlap with previous works on subsets of the domains we include (Janisch, 2024; Garg et al., 2019; Sharma et al., 2023).
@@ 4.1 Problem Selection
PyRDDLGym provides an archive of problem definitions written in RDDL. [^11] https://github.com/pyrddlgym-project/rddlrepository/commit/1a2d3b5 RDDL defines problems as a combination of a domain description file and an instance description file. The domain file defines state transition dynamics and object classes, as well as default values. The instance file declares object instances, ground values and parameters to construct a specific problem instance. RDDL is similar to the Planning Domain Definition Language (PDDL) in that it is a formal description language used to describe decision problems, but adds modeling features such as probabilities, continuous state variables and partial observability.
@@
We evaluated Vejde on the following problem domains from previous IPPCs: Elevators (2014), SysAdmin, Navigation, Traffic (2014), SkillTeaching (2014), AcademicAdvising (2014), CrossingTraffic (2014) and Tamarisk. The parenthesized year denotes the version we used where multiple are available. Each domain has ten problem instances with variations to initial values and the number of entities. All problems have a time horizon of 40 steps, after which the MDP is terminated, and all problems have the option for the agent to do nothing. We selected domains based on a set of exclusionary criteria, that filter out problems that define continuous actions, action literals with an arity greater than one, state-dependent conditions on actions and lastly problems that implement partial observability. For example, the IPPC 2014 problem Wildfire were excluded since it contains the action fluent put_out(x,y)put\_out(x,y), which has arity of two. Modifying domains specifications to include more problems is possible, but we considered it out of scope for this work. While we did not modify the domains, we edited all instance specifications that allow for concurrent agent actions to only permit one action per timestep. This includes subsets of instances in SkillTeaching, AcademicAdvising, Elevators and all instances in Traffic. Removing concurrent actions allows us to compare the GNN policy against the pyRDDLGym MLP policy on all instances, at the expense of slightly changing the problem premises. [^12] Without concurrent actions for Elevators, an agent has to prioritize moving one of two elevators at each time step rather than controlling both simultaneously. The filter conditions used to select problems are discussed further in Section 7.
@@
The included domains cover two general types of decision problems: SysAdmin, Elevators, SkillTeaching, Tamarisk and Traffic are reward maximization, or cost minimization, tasks. CrossingTraffic, Navigation and AcademicAdvising are goal-oriented, where the agent actions should lead to a given condition being fulfilled. Based on the arities of the action literals defined for the problems, we can also separate the eight problem domains into three categories. Ignoring the “do nothing” action, SysAdmin, Traffic and AcademicAdvising have only one unary action and no nullary actions. Elevators, SkillTeaching and Tamarisk have multiple unary actions and no nullary actions. CrossingTraffic, Navigation have four nullary actions that represent cardinal directions for the agent to move, but no unary actions. The maximum predicate arity in all domains is two. This fact was not part of our inclusion criteria, but came as a consequence of the selected problems. Vejde can be used with the different domains with no modifications other than adjusting the input and output sizes for each domain, meaning that we train one agent per domain. The input and output sizes are dependent on the number of predicate symbols and action symbols in the domain respectively.
@@ 4.2 Data Generation
A RDDL problem instance defines a constant number of objects. Since we are interested in problems where the total number of objects vary, or is unknown, we sample transitions evenly from multiple problem instances as if they constitute a single relational MDP. To verify that the policy is inductive, we divide the 10 problem instances of each domain into two sets by random selection and use one for training and the other for testing, as per common practice in supervised learning.
By the classification of Kirk et al. (2023) we are thus focusing on Independent and Identically Distributed (IID) generalization of the agent. The number of objects in the problem instances are not evenly distributed; to reduce result bias caused by the selection of instances for the training set, we sample five train/test splits and repeat experiments for each split, and calculate the final results as an average over them. For the other agents, Prost and the MLPs, the notion of a “test” set of problems does not work, as Prost always performs searches and the MLP agents are tied to particular problem instances.
@@ 4.3.1 Imitation Learning
We tested the capability of Vejde to encode a near-optimal policy by having agents mimic the actions of the planning algorithm Prost on the training set of problems. The resulting mimic policy was then run against the test set of problems. To collect training data for each domain, we ran Prost on every instance in the training set for ten episodes, each consisting of 40 state-action transitions with the recommended “IPC2014” configuration. This resulted in 2000 samples for each domain, each of which consists of a state and an action taken by Prost. The mimic agent was then trained through imitation learning, where the policy is optimized to simply maximize the probability of actions selected by Prost. This procedure was repeated five times, once for each train/test split. The results of this experiment can be found in Section 5.1.
@@ 4.3.2 Reinforcement Learning
As we do not wish to assume access to labeled data or an expert policy, the primary interest for us with this work is to find decision policies using RL. While we solely used PPO for RL, other policy gradient algorithms could be used in theory. In training the agent, we maximize the loss
$L=-L_{a}-c_{c}L_{c}+c_{h}H(\pi)$
through stochastic gradient ascent, where $L_{a}$ is the actor loss, $L_{c}$ the critic prediction loss ${\left(\tilde{v}-R_{t}\right)}^{2}$ and $H(\pi)$ the entropy of the action distribution given the state. The entropy term serves as a regularization parameter and encourages exploration. Each Vejde policy was trained with a total of 1 500 0001\,500\,000 samples from problem instances in the training set. All MLP policies were trained with 400 000400\,000 samples per instance with the same hyperparameters for all problems. Though both the MLP and Vejde agents are trained as stochastic policies, we evaluated them deterministically by picking the most probable action from the action distribution for a given state, rather than sampling the distribution. This choice was made to lower the variance of the results, but it ignores that the policy may assign nearly equal probabilities to some actions.
We used the same set of hyperparameters while training the GNN for all domains, which were chosen manually based on training time and returns observed in preliminary experiments on the training set. A rudimentary annealing scheme was used while training Vejde agents, where the learning rate was lowered by a factor of 10 for every 500 000500\,000 samples. The coefficient of the entropy term was also lowered after 500 000500\,000 samples, to guard against the policy converging to local minima early. The full set of hyperparameters is listed in Appendix A.
A source of difficulty in assigning a single set of hyperparameters was that the return values of the different domains, and even instances within the same domain, can have vastly different magnitudes. To improve the robustness of the optimization, we used two tricks from Dreamer (Hafner et al., 2025). The first is to scale $R_{t}$ and the value estimate $\tilde{v}$ by the symlog function, and the second is to scale the advantage $A_{t}$ by an exponential moving average of the return value range. The advantage $A_{t}$ was calculated using generalized advantage estimation (GAE(Schulman et al., 2016), as is common in implementations of PPO. The results of this experiment can be found in Section 5.2.
@@ 5 Results
Our main metric of success is the average return the agents receive on the problem instances in the test set, which the Vejde agents were not trained on. For the agents that are not inductive, i.e. Prost and the MLP agents, we only count scores on problems that are in the test set, averaged over the respective train/test splits.
Each RDDL domain uses rewards with different scales and magnitudes, which makes comparing results across domains, or instances, difficult. We thus present our results as scores normalized to the range $[0,1]$. Returns are normalized according to the method used in IPPC 2014 (Vallati et al., 2015): For each instance, a lower bound return, $R_{low}$, is the maximum average return received from taking random actions or doing nothing. A maximum return, $R_{\max}$, for an instance is the highest return obtained among the evaluated methods. The score of an agent on a given instance is then $\nicefrac{{\max(R-R_{low},0)}}{{(R_{\max}-R_{base})}}$, where $R$ is the average return over a given number of episodes. Thus, a score of 0 means that an agent performs worse or equal to acting at random or doing nothing, and 1 represents always having the highest score among the compared agents. We emphasize that a score of 1 does not imply that the agent follows the optimal policy of the MDP, which we do not have access to for comparison. Raw return values for each domain, instance and agent can be obtained from our GitHub repository. [^13] https://github.com/kasanari/vejde-rddl-evalAll experiments were done on a machine equipped with 32 GB of RAM, an Intel Xeon Silver CPU with 24 cores and an NVIDIA Quadro RTX 4000 GPU.
@@ 5.1 Imitation Learning
Running Prost on the 90 problems for ten episodes each to collect training data took approximately 6 hours. We then ran Prost for 100 episodes on all 90 problems for scores to use in evaluation, which took roughly two days. Each mimic policy was trained for 1000 epochs over the training data, which was fed to the agent in single batches of 2000 samples. Training mimic policies for all domains took approximately 30 minutes, repeated five times with different train-test splits for a total of 2.5 hours. The mimic policy was then tested on all problem instances for 100 episodes. To summarize the differences between the agents, we performed a permutation test to compare average mimic scores on train and test instances, as well as Prost scores. [^14] We used https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.permutation_test.html to run the test. The test calculates a distribution of a selected metric, in our case the difference in means, under the hypothesis that the compared sample distributions are equal. It does so by pooling all the samples and calculating the metric for permutations of the samples for a given number of repetitions. If the observed metric is unlikely given the resulting distribution of metrics, then the null hypothesis of the samples have equal means can be rejected. We observed a difference of 0.01 between the agent scores on test problems compared to train problems, which according to the null distribution has a probability of occurring by 83%. We thus have more confidence in the means being equal than not according to the test. The differences between Prost and both the training and test sets are significantly larger, with a difference of -0.26 between Prost and the Vejde scores on the test set. The differences in mean with the $p$-values from the permutation test is shown in Table 1, and a plot of the null distribution from the test is shown in Appendix B.
@@ 5.2 Reinforcement Learning
Training Vejde policies for all domains using reinforcement learning took approximately 5 hours, which repeated five times took a total of 35 hours. Training one MLP policy for each of the 90 problem instances took approximately 24 hours. After training, all policies were executed on each of the problem instances for 100 episodes, and the average return per instance was recorded. We used the same evaluation data for Prost as was used in the imitation learning experiment. To create a summary score for over all domains, we performed a permutation test with $10^{6}$ samples to compare the test problem scores of the Vejde agents with the scores of the MLP policies and Prost on the same problems, with the null hypothesis that the agents have equal average scores. If we assume that we can reject the null hypothesis with $p<0.05$, we can not do so for the difference between the mean scores of the MLP and Vejde agents, but we can for the difference between Prost and the other agents. There was thus, in average, not a significant statistical difference between the average scores of the Vejde agents and the MLP agents, which we consider good given that the Vejde agents were not trained on the considered instances. Both RL agents have significantly worse scores than Prost in average. The statistics of the test are shown in Table 3, and we have included a plot of the resulting null distribution from running the test in Appendix B.

### v2

@@ 3.5 Value Estimate
where $\widetilde{Q}(u,a)=h_{u}\cdot W_{a};\ W\in\mathbb{R}^{D\times|A|},\ h_{u}\in H.$ This formulation can be regarded as dividing a single step in the MDP into two sub-steps, where the first consists of the agent choosing a predicate symbol. If we were to extend the number of object arguments for actions, the number of sub-steps would increase to account for the additional choices.
@@ 4 Evaluation
We evaluate Vejde using a set of problems defined with the Relational Dynamic Influence Diagram Language (RDDL) and executed in the Python library PyRDDLGym [^9] https://github.com/pyrddlgym-project/pyRDDLGym/commit/f7dd1dd (taitler2024pyrddlgymrddlgymenvironments). RDDL is a description language that can be used for specifying Markov decision problems as dynamic Bayesian networks (sanner2010relational). We test the ability of Vejde to model inductive decision policies in two ways: supervised learning from examples and reinforcement learning.
@@
We compare Vejde policies with policies parametrized using MLPs, and a planning algorithm named Prost (keller-helmert-icaps2013), the winner of the probabilistic track of the 2014 International Probabilistic Planning Competition (IPPC(Vallati_Chrpa_Grześ_McCluskey_Roberts_Sanner_Editor_2015). Prost uses a tree-search algorithm framework to find an optimal action for each state based on a series of simulations initiated from the state. As such, for each state Prost runs thousands of trials to construct the search-tree. We think it should be noted that this is a fundamentally different problem-solving approach than the reinforcement learning method used to train the Vejde and MLP policy functions. With RL, a simulator is only used during training to find the parameters of the policies, whereas Prost always needs access to a simulator of the problem in order to perform searches.
We use an existing interface for Prost to PyRDDLGym for all experiments. [^10] https://github.com/pyrddlgym-project/pyRDDLGym-prost/commit/248d5d2 For the MLP policies, we used the implementation included in the pyRDDLGym set of libraries, which in turn is based on Stable Baselines 3.111111https://github.com/pyrddlgym-project/pyRDDLGym-rl/commit/9714392 In the MLP implementation, the state is represented as a vector over all possible facts, the length of which is dependent on the language and number of objects in the problem instance. This means that the total parameter count of the MLPs input and output layers varies between instances and domains, but all use two hidden layers with 64 latent parameters each. The MLP policies are transductive, and can not be used for multiple problem instances due to the varying input and output shapes. We thus train one MLP agent for each domain and instance combination.
Including an inductive agent from previous works would have been useful, but we were unable to find a single method that cover the set of problem domains as we are evaluating on. We should note, however, that there is overlap with previous works on subsets of the domains we include (janisch2024applications; DBLP:conf/aips/GargBM19; DBLP:conf/uai/SharmaAMS23). We did evaluate an alternative encoding method, however, based on graph attention. The main difference between this approach and the message passing method described in Section 3.3 is that the graph attention method incorporates information from the entire graph in a single step of message passing. We ultimately found that the graph attention method did not perform significantly better than the message passing method, at a significantly higher computational cost. An extended description of the graph attention method, together with evaluation results, is shown in Appendix H.
@@ 4.1 Problem Selection
RDDL defines problems as a combination of a domain description file and an instance description file. The domain file primarily defines state transition dynamics and object classes, whereas instance files declares object instances, ground values and parameters to instantiate problems in the domain. Vejde can be used with the different domains with no modifications other than adjusting the input and output sizes based on each domain description, meaning that we train one agent per domain. RDDL is similar to the Planning Domain Definition Language (PDDL) in that it is a formal description language used to describe decision problems, but adds modeling features such as probabilities, continuous state variables and partial observability.
@@
PyRDDLGym hosts a repository of RDDL problem definitions [^12] https://github.com/pyrddlgym-project/rddlrepository/commit/1a2d3b5. We evaluated Vejde on the following problem domains from previous IPPCs: Elevators (2014), SysAdmin, Navigation, Traffic (2014), SkillTeaching (2014), AcademicAdvising (2014), CrossingTraffic (2014) and Tamarisk [^13] The parenthesized year denotes the version used where multiple are available.. Each problem domain has ten instances which vary initial values and the number of entities. All problems have a time horizon of 40 steps, after which the MDP is terminated, and all problems have the option for the agent to do nothing. We selected domains based on a set of exclusionary criteria that filter out problems that define continuous actions, action literals with an arity greater than one, state-dependent conditions on actions and lastly problems that implement partial observability. For example, the IPPC 2014 problem Wildfire were excluded since it contains the action fluent put_out(x,y)put\_out(x,y), which has arity of two. Modifying domains specifications to include more problems is possible, but we considered it out of scope for this work. While we did not modify the domains, we edited all instance specifications that allow for concurrent agent actions to only permit one action per timestep. This includes subsets of instances in SkillTeaching, AcademicAdvising, Elevators and all instances in Traffic. Removing concurrent actions allows us to compare the GNN policy against the pyRDDLGym MLP policy on all instances, at the expense of slightly changing the problem premises. [^14] Without concurrent actions for Elevators, an agent has to prioritize moving one of two elevators at each time step rather than controlling both simultaneously. The conditions used to select problems are discussed further in Appendix D.
@@
The included domains cover two general types of decision problems: SysAdmin, Elevators, SkillTeaching, Tamarisk and Traffic are reward maximization, or cost minimization, tasks. CrossingTraffic, Navigation and AcademicAdvising are goal-oriented, where the agent actions should lead to a given condition being fulfilled. Based on the arities of the action literals defined for the problems, we can also separate the eight problem domains into three categories. Ignoring the “do nothing” action, SysAdmin, Traffic and AcademicAdvising have only one unary action and no nullary actions. Elevators, SkillTeaching and Tamarisk have multiple unary actions and no nullary actions. CrossingTraffic, Navigation have four nullary actions that represent cardinal directions for the agent to move, but no unary actions. The maximum predicate arity in all domains is two. This fact was not part of our inclusion criteria, but came as a consequence of the selected problems.
@@ 4.2 Data Generation
A RDDL problem instance defines a constant number of objects. As we are interested in use-cases where the total number of objects vary, or is unknown, we sample transitions evenly from multiple problem instances as if they constitute a single relational MDP. To test that the policy generalizes, we divide the 10 problem instances of each domain into two sets by random selection and use one for training and the other for testing, as per common practice in supervised learning. The object counts are not evenly distributed, so to reduce bias we sample five train/test splits and repeat experiments for each split, and calculate the final results as an average over them. For the other agents, Prost and the MLPs, considering a “test” set of problems is not possible, as Prost always performs searches and each MLP agent is tied to a problem instance.
@@ 4.3.1 Imitation Learning
We tested the capability of Vejde to encode a near-optimal policy by having agents mimic the actions of the planning algorithm Prost on the training set of problems. The resulting mimic policy was then scored on the test set of problems. To collect training data for each domain, we ran Prost on every instance in the training set for ten episodes, each consisting of 40 state-action transitions with the recommended “IPC2014” configuration. This resulted in 2000 samples for each domain, each of which consists of a state and an action taken by Prost. The mimic agent was then trained through imitation learning, where the policy is optimized to simply maximize the probability of actions selected by Prost. This procedure was repeated five times, once for each train/test split. The results of this experiment can be found in Section 5.1.
@@ 4.3.2 Reinforcement Learning
As we do not wish to assume access to labeled data or an expert policy, the primary interest for us with this work is to find decision policies using RL. While we solely used PPO for RL, other policy gradient algorithms could be used in theory. In training the agent, we maximize the loss $L=-L_{a}-c_{c}L_{c}+c_{h}H(\pi)$ through stochastic gradient ascent, where $L_{a}$ is the actor loss, $L_{c}$ the critic prediction loss ${\left(\tilde{v}-R_{t}\right)}^{2}$ and $H(\pi)$ the entropy of the action distribution given the state. The entropy term serves as a regularization parameter and encourages exploration. Each Vejde policy was trained with a total of 1 500 0001\,500\,000 samples from problem instances in the training set. All MLP policies were trained with 400 000400\,000 samples per instance with the same hyperparameters for all problems. Though both the MLP and Vejde agents are trained as stochastic policies, we evaluated them deterministically by picking the most probable action from the action distribution for a given state, rather than sampling the distribution. This choice was made to lower the variance of the results, but it ignores that the policy may assign nearly equal probabilities to some actions. The results of this experiment can be found in Section 5.2.
@@ 5 Results
Our primary evaluation metric is the average return for agents on problem instances in the test set, which Vejde agents were not trained on. For transductive agents, i.e. Prost and the MLP agents, we only count scores on problems that are in the test set, averaged over the respective train/test splits.
Each RDDL domain uses rewards with different scales and magnitudes. We thus present returns normalized to the range $[0,1]$, calculated using the method from IPPC 2014 (Vallati_Chrpa_Grześ_McCluskey_Roberts_Sanner_Editor_2015). Raw return values for each domain, instance and agent can be obtained from our GitHub repository [^15] https://github.com/kasanari/vejde-rddl-eval. All experiments were done on a machine equipped with 32 GB of RAM, an Intel Xeon Silver CPU with 24 cores and an NVIDIA Quadro RTX 4000 GPU.
@@ 5.1 Imitation Learning
Running Prost on the 90 problems for ten episodes each to collect training data took approximately 6 hours. We then ran Prost for 100 episodes on all 90 problems for scores to use in evaluation, which took roughly two days. Each mimic policy was trained for 1000 epochs over the training data, which was fed to the agent in single batches of 2000 samples. Training mimic policies for all domains took approximately 30 minutes, repeated five times with different train-test splits for a total of 2.5 hours. The mimic policy was then tested on all problem instances for 100 episodes. To summarize the differences between the agents, we performed a permutation test to compare average mimic scores on train and test instances, as well as Prost scores. [^16] We used https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.permutation_test.html to run the test. The test calculates a distribution of a selected metric, in our case the difference in means, under the hypothesis that the compared sample distributions are equal. It does so by pooling all the samples and calculating the metric for permutations of the samples for a given number of repetitions. If the observed metric is unlikely given the resulting distribution of metrics, then the null hypothesis of the samples have equal means can be rejected. We observed a difference of 0.01 between the agent scores on test problems compared to train problems, which according to the null distribution has a probability of occurring by 83%. We thus have more confidence in the means being equal than not according to the test. The differences between Prost and both the training and test sets are significantly larger, with a difference of -0.26 between Prost and the Vejde scores on the test set. The differences in mean with the $p$-values from the permutation test is shown in Table 1, and a plot of the null distribution from the test is shown in Appendix F.
@@ 5.2 Reinforcement Learning
Training Vejde policies for all domains using reinforcement learning took approximately 5 hours, which repeated five times took a total of 35 hours. Training one MLP policy for each of the 90 problem instances took approximately 24 hours. After training, all policies were executed on each of the problem instances for 100 episodes, and the average return per instance was recorded. We used the same evaluation data for Prost as was used in the imitation learning experiment. To create a summary score for over all domains, we performed a permutation test with $10^{6}$ samples to compare the test problem scores of the Vejde agents with the scores of the MLP policies and Prost on the same problems, with the null hypothesis that the agents have equal average scores. If we assume that we can reject the null hypothesis with $p<0.05$, we can not do so for the difference between the mean scores of the MLP and Vejde agents, but we can for the difference between Prost and the other agents. There was thus, in average, not a significant statistical difference between the average scores of the Vejde agents and the MLP agents, which we consider good given that the Vejde agents were not trained on the considered instances. Both RL agents have significantly worse scores than Prost in average. The statistics of the test are shown in Table 3, and we have included a plot of the resulting null distribution from running the test in Appendix F.